Show that sinØ = cos(90deg - Ø)
Answer by lwsshak3(11628)

Show that sinØ = cos(90deg -Ø)
Use cos addition formula:
cos(90-Ø)=cos(90)cos(Ø)+sin(90)*sinØ=0*cos(Ø)+1*sin(Ø)=sinØ